OpenStudy (anonymous):

(__)^2= (csc x-1)(csc x+1) solve for the blank

OpenStudy (anonymous):

would it be cscx^2

OpenStudy (btaylor):

it would be \[\csc^{2} x\]

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

(csc x-1)(csc x+1) = \[\csc^{2}(x) - 1 = (blank)^{2}\] so \[\sqrt{\csc2(x)−1} = blank\]

OpenStudy (anonymous):

Nolastudent-- you're way makes sense, but is there a way to simplify is more?

OpenStudy (anonymous):

yes, sorry, trig identity: csc^2(x)-1 = cot^2(x)

OpenStudy (anonymous):

*your*. Sorry... :)

OpenStudy (anonymous):

so square root of cotangent squared, which is cotangent.

OpenStudy (anonymous):

That's okay, we're in the math section XD

OpenStudy (anonymous):

Okay, that makes more sense. I can never simplify mine enough... :)

OpenStudy (anonymous):

LOL, thanks for the help gandalfwiz!

OpenStudy (anonymous):

So the final answer is cotx?

OpenStudy (anonymous):

yep